### v1

# Coarse spaces using extended generalized eigenproblems for heterogeneous Helmholtz problems

11institutetext: Frédéric Nataf 22institutetext: Sorbonne Université, Université Paris Cité, CNRS, INRIA, Laboratoire Jacques-Louis Lions, LJLL, EPC ALPINES, F-75005 Paris, France, 22email: frederic.nataf@sorbonne-universite.fr 33institutetext: Emile Parolin 44institutetext: Sorbonne Université, Université Paris Cité, CNRS, INRIA, Laboratoire Jacques-Louis Lions, LJLL, EPC ALPINES, F-75005 Paris, France 44email: emile.parolin@inria.fr
Coarse spaces using extended generalized eigenproblems for heterogeneous Helmholtz problems
[…]
1 Optimized Restricted Additive Schwarz method
Model problem. Let $d\in\{1,2,3\}$ and $\Omega\subset\mathbb{R}^{d}$ be an open and bounded domain. We consider the heterogeneous Helmholtz equation for $\omega>0$
$\text{Find}\ u\in H^{1}(\Omega)\ \text{such that:}\qquad\begin{cases}-\nabla\cdot(\mu\nabla u)-\omega^{2}\nu u=f,&\text{in}\ \Omega,\\ \mu\partial_{\mathbf{n}}u-\imath\omega u=g,&\text{on}\ \partial\Omega,\end{cases}$ (1)
where $\mu$ and $\nu$ are two variable but positive and bounded coefficients, $f\in L^{2}(\Omega)$ and $g\in L^{2}(\partial\Omega)$ model the source terms, and $\mathbf{n}$ is the outward unit normal vector to the boundary $\partial\Omega$.
The model problem (1) takes the following variational form
$\text{Find}\ u\in H^{1}(\Omega)\ \text{such that:}\qquad a_{\Omega}(u,u^{t})=l_{\Omega}(u^{t}),\quad\forall u^{t}\in H^{1}(\Omega),$ (2)
where we introduced, for any $D\subset\Omega$ and any $u,u^{t}\in H^{1}(D)$:
$a_{D}(u,u^{t})$ $:=(\mu\nabla u,\nabla u^{t})_{L^{2}(D)^{d}}-\omega^{2}(\nu u,u^{t})_{L^{2}(D)}-\imath\omega(u,u^{t})_{L^{2}(\partial D\cap\partial\Omega)},$ (3)
$l_{D}(u^{t})$ $:=(f,u^{t})_{L^{2}(D)}+(g,u^{t})_{L^{2}(\partial D\cap\partial\Omega)}.$ (4)
The above model problem (2) is well-posed using standard arguments based on the Fredholm alternative and a unique continuation principle.
One-level domain decomposition algorithm. We introduce a decomposition of $\Omega$ into $J\in\mathbb{N}$, $J>0$, overlapping open subdomains $\Omega_{j}$, $j=1,\ldots,J$, associated to a partition-of-unity function $\chi_{j}\in W^{1,\infty}(\Omega_{j})$ such that
$\overline{\Omega}=\bigcup_{j}\overline{\Omega_{j}},\qquad\chi_{j}\geq 0,\ \chi_{j}|_{\partial\Omega_{j}\setminus\partial\Omega}=0,\qquad\sum_{j}E_{j}\chi_{j}u|_{\Omega_{j}}=u,\ \forall u\in H^{1}(\Omega),$ (5)
where $E_{j}$ is an extension-by-continuity operator. More precisely, we require that $E_{j}:H^{1}(\Omega_{j})\to H^{1}({\Omega})$ continuously and such that $\operatorname{supp}(E_{j}u_{j})\subset\widetilde{\Omega}_{j}$ for any $u_{j}\in H^{1}(\Omega_{j})$ and for some extended open subdomain $\widetilde{\Omega}_{j}$ such that ${\Omega}_{j}\subsetneq\widetilde{\Omega}_{j}\subset\Omega$.
Starting from an initial guess $u^{0}_{\mathrm{I}}\in H^{1}(\Omega)$, the one-level (the index I stands for one-level) Optimized Restricted Additive Schwarz (ORAS) algorithm defines the sequence
$u^{n+1}_{\mathrm{I}}:=u^{n}_{\mathrm{I}}+\sum_{j}E_{j}\chi_{j}v_{j}^{n+1},\qquad n\geq 0,$ (6)
where $v_{j}^{n+1}$ is solution of: Find $v_{j}^{n+1}\in H^{1}(\Omega_{j})$ such that:
$b_{\Omega_{j}}(v_{j}^{n+1},v_{j}^{t})=l_{\widetilde{\Omega}_{j}}(E_{j}v_{j}^{t})-a_{\widetilde{\Omega}_{j}}(u^{n}_{\mathrm{I}}|_{\widetilde{\Omega}_{j}},E_{j}v_{j}^{t}),\qquad\forall v^{t}_{j}\in H^{1}(\Omega_{j}),$ (7)
where we introduced, for any $D\subset\Omega$ and any $v,v^{t}\in H^{1}(D)$:
$b_{D}(v,v^{t}):=a_{D}(v,v^{t})-\imath\omega(v,v^{t})_{L^{2}(\partial D\setminus\partial\Omega)}.$ (8)
By similar arguments as for the model problem, the local problem (7) is well-posed.
One-level error analysis. For $n\geq 0$, the one-level error is defined as
$e^{n+1}_{\mathrm{I}}:=u-u^{n+1}_{\mathrm{I}}=e^{n}_{\mathrm{I}}-\sum_{j}E_{j}\chi_{j}v_{j}^{n+1}=\sum_{j}E_{j}\chi_{j}(e^{n}_{\mathrm{I}}|_{\Omega_{j}}-v_{j}^{n+1}).$ (9)
From (7), we have,
$b_{\Omega_{j}}(e^{n}_{\mathrm{I}}|_{\Omega_{j}}-v_{j}^{n+1},u_{j}^{t})=b_{\Omega_{j}}(e^{n}_{\mathrm{I}}|_{\Omega_{j}},v_{j}^{t})-a_{\widetilde{\Omega}_{j}}(e^{n}_{\mathrm{I}}|_{\widetilde{\Omega}_{j}},E_{j}v_{j}^{t}),\ \forall v_{j}^{t}\in H^{1}(\Omega_{j}),$ (10)
so that
$e^{n+1}_{\mathrm{I}}=\sum_{j}E_{j}\chi_{j}R_{j}e^{n}_{\mathrm{I}}|_{\widetilde{\Omega}_{j}},$ (11)
where $R_{j}:H^{1}(\widetilde{\Omega}_{j})\to H^{1}(\Omega_{j})$ is defined for any $\widetilde{v}_{j}\in H^{1}(\widetilde{\Omega}_{j})$ as
$b_{\Omega_{j}}(R_{j}\widetilde{v}_{j},v_{j}^{t})=b_{\Omega_{j}}(\widetilde{v}_{j}|_{\Omega_{j}},v_{j}^{t})-a_{\widetilde{\Omega}_{j}}(\widetilde{v}_{j},E_{j}v_{j}^{t}),\ \forall v_{j}^{t}\in H^{1}(\Omega_{j}).$ (12)
The operator $R_{j}$ is well-defined since the local problems (12) are well-posed.
We measure the error in the energy norm given by the following sesquilinear form: for any $D\subset\Omega$ and any $u,u^{t}\in H^{1}(D)$, let
$c_{D}(u,u^{t}):=(\mu\nabla u,\nabla u^{t})_{L^{2}(D)^{d}}+\omega^{2}(\nu u,u^{t})_{L^{2}(D)},\qquad\|u\|_{D}^{2}:=c_{D}(u,u).$ (13)
The error can be estimated as follows
$\|e^{n+1}_{\mathrm{I}}\|_{\Omega}^{2}\leq k_{0}\sum_{j}\|\chi_{j}R_{j}e^{n}_{\mathrm{I}}|_{\widetilde{\Omega}_{j}}\|_{\Omega_{j}}^{2}\leq k_{0}\xi\sum_{j}\|e^{n}_{\mathrm{I}}|_{\widetilde{\Omega}_{j}}\|_{\widetilde{\Omega}_{j}}^{2}\leq k_{0}k_{1}\xi\|e^{n}_{\mathrm{I}}\|_{\Omega}^{2},$ (14)
where we introduced
$k_{0}:=\!\!\sup_{\begin{subarray}{c}(u_{j})_{j}\\ 0\neq u_{j}\in H^{1}_{0}(\Omega_{j})\end{subarray}}\!\!\frac{\|\sum_{j}E_{j}u_{j}\|_{\Omega}^{2}}{\sum_{j}\|u_{j}\|_{\Omega_{j}}^{2}},\qquad k_{1}:=\!\!\sup_{0\neq u\in H^{1}(\Omega)}\!\!\frac{\sum_{j}\|u|_{\widetilde{\Omega}_{j}}\|_{\widetilde{\Omega}_{j}}^{2}}{\|u\|_{\Omega}^{2}},$ (15)
$\xi:=\sup_{j}\xi_{j},\qquad\xi_{j}:=\!\!\sup_{0\neq\widetilde{u}_{j}\in H^{1}(\widetilde{\Omega}_{j})}\!\!\frac{\|\chi_{j}R_{j}\widetilde{u}_{j}\|_{\Omega_{j}}^{2}}{\|\widetilde{u}_{j}\|_{\widetilde{\Omega}_{j}}^{2}}.$ (16)
The estimate (14) shows that a sufficient condition to have convergence of the algorithm is that $k_{0}k_{1}\xi<1$, which is unlikely to be satisfied. The two constants $k_{0}$ and $k_{1}$ can be bounded uniformly of the parameters $\mu,\nu,\omega$ of the problem and only depend respectively on the maximum number of neighbors for a subdomain (see e.g. (Dolean2015, , Lem. 7.9)) and the maximum number of subdomain intersection multiplicity (see e.g. (Dolean2015, , Lem. 7.13)). For reasonable partitions, both quantities can be bounded uniformly of the number of subdomains. On the other hand, $\xi$ depends on the parameters of the problem. In the next section, we propose the construction of a coarse space and associated two-level domain decomposition algorithm that allows to control each $\xi_{j}$ independently, hence $\xi$.
2 Two-level ORAS method
Local generalized eigenproblem The definition of $\xi_{j}$ suggests considering the following local generalized eigenproblem: Find $(\widetilde{u}_{j},\lambda_{j})\in H^{1}(\widetilde{\Omega}_{j})\times\mathbb{R}$ such that:
$c_{\Omega_{j}}(\chi_{j}R_{j}\widetilde{u}_{j},\chi_{j}R_{j}\widetilde{u}_{j}^{t})=\lambda_{j}c_{\widetilde{\Omega}_{j}}(\widetilde{u}_{j},\widetilde{u}_{j}^{t}),\qquad\forall\widetilde{u}_{j}^{t}\in H^{1}(\widetilde{\Omega}_{j}).$ (17)
This problem is equivalent to the following saddle-point formulation:
Find $(\widetilde{u}_{j},v_{j},\sigma_{j},\lambda_{j})\in H^{1}(\widetilde{\Omega}_{j})\times H^{1}(\Omega_{j})\times H^{1}(\Omega_{j})\times\mathbb{R}$ such that:
$\begin{cases}c_{\Omega_{j}}(\chi_{j}v_{j},\chi_{j}v_{j}^{t})-b_{\Omega_{j}}(v_{j}^{t},\sigma_{j})=0,&\quad\forall v_{j}^{t}\in H^{1}(\Omega_{j}),\\ -b_{\Omega_{j}}(v_{j},\sigma_{j}^{t})+b_{\Omega_{j}}(\widetilde{u}_{j}|_{\Omega_{j}},\sigma_{j}^{t})-a_{\widetilde{\Omega}_{j}}(\widetilde{u}_{j},E_{j}\sigma_{j}^{t})=0,&\quad\forall\sigma_{j}^{t}\in H^{1}(\Omega_{j}),\\ b_{\Omega_{j}}(\widetilde{u}_{j}^{t}|_{\Omega_{j}},\sigma_{j})-a_{\widetilde{\Omega}_{j}}(\widetilde{u}_{j}^{t},E_{j}\sigma_{j})=\lambda_{j}c_{\widetilde{\Omega}_{j}}(\widetilde{u}_{j},\widetilde{u}_{j}^{t}),&\quad\forall\widetilde{u}_{j}^{t}\in H^{1}(\widetilde{\Omega}_{j}).\end{cases}$ (18)
The following result shows that each $\xi_{j}$ (hence $\xi$) is finite and that the eigenvalues $\lambda_{j}$ solutions to (17) converge to 0.
Proposition 1
The operator $K_{j}:\widetilde{u}_{j}\in H^{1}(\widetilde{\Omega}_{j})\mapsto\chi_{j}R_{j}\widetilde{u}_{j}\in H^{1}(\Omega_{j})$ is compact.
Proof
The result is based on a Caccioppoli inequality which we establish next (see also (Ma2023, , Lem. 2.7)). From (12) we have, for any $\widetilde{u}_{j}\in H^{1}(\widetilde{\Omega}_{j})$,
$b_{\Omega_{j}}(R_{j}\widetilde{u}_{j},\chi_{j}^{2}R_{j}\widetilde{u}_{j})=0,$ (19)
which implies (adding (19) and its conjugate) that
$(\mu\chi_{j}\nabla R_{j}\widetilde{u}_{j},\chi_{j}\nabla R_{j}\widetilde{u}_{j})_{L^{2}(\Omega_{j})^{d}}+(\mu R_{j}\widetilde{u}_{j}\nabla\chi_{j},\chi_{j}\nabla R_{j}\widetilde{u}_{j})_{L^{2}(\Omega_{j})^{d}}$
$\quad+(\mu\chi_{j}\nabla R_{j}\widetilde{u}_{j},R_{j}\widetilde{u}_{j}\nabla\chi_{j})_{L^{2}(\Omega_{j})^{d}}-\omega^{2}(\nu R_{j}\widetilde{u}_{j},\chi_{j}^{2}R_{j}\widetilde{u}_{j})_{L^{2}(\Omega_{j})}=0.$
On the other hand, we have, for any $\widetilde{u}_{j}\in H^{1}(\widetilde{\Omega}_{j})$,
$\|\chi_{j}R_{j}\widetilde{u}_{j}\|_{\Omega_{j}}^{2}=c_{\Omega_{j}}(\chi_{j}R_{j}\widetilde{u}_{j},\chi_{j}R_{j}\widetilde{u}_{j})=(\mu R_{j}\widetilde{u}_{j}\nabla\chi_{j},R_{j}\widetilde{u}_{j}\nabla\chi_{j})_{L^{2}(\Omega_{j})^{d}}$
$\quad+(\mu\chi_{j}\nabla R_{j}\widetilde{u}_{j},\chi_{j}\nabla R_{j}\widetilde{u}_{j})_{L^{2}(\Omega_{j})^{d}}+(\mu R_{j}\widetilde{u}_{j}\nabla\chi_{j},\chi_{j}\nabla R_{j}\widetilde{u}_{j})_{L^{2}(\Omega_{j})^{d}}$
$\quad+(\mu\chi_{j}\nabla R_{j}\widetilde{u}_{j},R_{j}\widetilde{u}_{j}\nabla\chi_{j})_{L^{2}(\Omega_{j})^{d}}+\omega^{2}(\nu\chi_{j}R_{j}\widetilde{u}_{j},\chi_{j}R_{j}\widetilde{u}_{j})_{L^{2}(\Omega_{j})}$
$=(\mu R_{j}\widetilde{u}_{j}\nabla\chi_{j},R_{j}\widetilde{u}_{j}\nabla\chi_{j})_{L^{2}(\Omega_{j})^{d}}+2\omega^{2}(\nu R_{j}\widetilde{u}_{j},\chi_{j}^{2}R_{j}\widetilde{u}_{j})_{L^{2}(\Omega_{j})}\leq C\|R_{j}\widetilde{u}_{j}\|^{2}_{L^{2}(\Omega_{j})},$
for some constant $C>0$ depending on $\omega$, upper bounds on the coefficients $\mu,\nu$ and the partition-of-unity $\chi_{j}$. The result then follows by application of the Rellich–Kondrachov theorem.
Coarse space. For a user-defined parameter $\tau>0$, we introduce the coarse space
[…]
The coarse space $U_{\tau}$ is made of local Helmholtz solutions $R_{j}\widetilde{u}_{j}$ that are extended-by-zero globally thanks to $E_{j}$ after being multiplied by the partition-of-unity function $\chi_{j}$. Indeed, for $\widetilde{u}_{j}$ solution of (17), we have $R_{j}\widetilde{u}_{j}=v_{j}$ solution of (18) which satisfies (testing with $\sigma_{j}^{t}\in H^{1}_{0}(\Omega_{j})$ in (18))
$(\mu\nabla v_{j},\nabla v_{j}^{t})_{L^{2}(\Omega_{j})^{d}}-\omega^{2}(\nu v_{j},v_{j}^{t})_{L^{2}(\Omega_{j})}=0,\qquad\forall v_{j}^{t}\in H^{1}_{0}(\Omega_{j}).$ (21)
Moreover, thanks to the partition-of-unity, the coarse space is conforming, namely $U_{\tau}\subset H^{1}(\Omega)$. Finally, Proposition 1 ensures that $U_{\tau}$ is finite dimensional.
Two-level domain decomposition algorithm. Starting from an initial guess $u^{0}_{\mathrm{II}}\in H^{1}(\Omega)$, we define a sequence of solutions $u^{n}_{\mathrm{II}}\in H^{1}(\Omega)$ (the index II stands for two-level),
$u^{n+1}_{\mathrm{II}}:=u^{n}_{\mathrm{II}}+\sum_{j}E_{j}\chi_{j}w_{j}^{n+1}+w_{0}^{n+1},\qquad n\geq 0,$ (22)
where $w_{j}^{n+1}$ is solution of: Find $w_{j}^{n+1}\in H^{1}(\Omega_{j})$ such that:
$b_{\Omega_{j}}(w_{j}^{n+1},w_{j}^{t})=l_{\widetilde{\Omega}_{j}}(E_{j}w_{j}^{t})-a_{\widetilde{\Omega}_{j}}(u^{n}_{\mathrm{II}}|_{\widetilde{\Omega}_{j}},E_{j}w_{j}^{t}),\qquad\forall w^{t}_{j}\in H^{1}(\Omega_{j}),$ (23)
and $w_{0}^{n+1}$ is solution of: Find $w_{0}^{n+1}\in U_{\tau}$ such that:
$a_{\Omega}(w_{0}^{n+1},w_{0}^{t})=l_{\Omega}(w_{0}^{t})-a_{\Omega}(u^{n}_{\mathrm{II}}+\sum_{j}E_{j}\chi_{j}w_{j}^{n+1},w_{0}^{t}),\quad\forall w_{0}^{t}\in U_{\tau}.$ (24)
Assumption 1
For any anti-linear form $L$ on $H^{1}(\Omega)$, consider the two problems
[…]
In the following, we assume that the coarse problem (26) is well-posed and that there exists $\sigma>0$, independent of $\tau$ and $L$, such that the following stability bound holds
$\|u-u_{0}\|_{\Omega}\leq\sigma\|u\|_{\Omega}.$ (27)
Assumption 1 implies that the problem (24) is well-posed. The validity of Assumption 1 can be argued for large enough coarse spaces $U_{\tau}$ (i.e. taking the user-controlled parameter $\tau$ sufficiently small). For instance, well-posedness of the coarse problem and a quasi-optimality estimate implying (27) can be obtained using the so-called Aubin–Nitsche duality trick and Schatz argument, see (Galkowski2025, , Lemma 5.1).
Two-level error analysis. For $n\geq 0$, the two-level error is defined as
$e^{n+1}_{\mathrm{II}}:=u-u^{n+1}_{\mathrm{II}}=u-u^{n}_{\mathrm{II}}-\sum_{j}E_{j}\chi_{j}w_{j}^{n+1}-w_{0}^{n+1}.$ (28)
From (23), we have,
$b_{\Omega_{j}}(e^{n}_{\mathrm{II}}|_{\Omega_{j}}-w_{j}^{n+1},w_{j}^{t})=b_{\Omega_{j}}(e^{n}_{\mathrm{II}}|_{\Omega_{j}},w_{j}^{t})-a_{\widetilde{\Omega}_{j}}(e^{n}_{\mathrm{II}}|_{\widetilde{\Omega}_{j}},E_{j}w_{j}^{t}),\ \forall w_{j}^{t}\in H^{1}(\Omega_{j}),$ (29)
and from (24), we have
$a_{\Omega}(w_{0}^{n+1},w_{0}^{t})=a_{\Omega}(\sum_{j}E_{j}\chi_{j}(e^{n}_{\mathrm{II}}|_{\Omega_{j}}-w_{j}^{n+1}),w_{0}^{t}),\quad\forall w_{0}^{t}\in U_{\tau},$ (30)
so that
$e^{n+1}_{\mathrm{II}}=(\mathrm{I}-P_{0})\sum_{j}E_{j}\chi_{j}R_{j}e^{n}_{\mathrm{II}}|_{\widetilde{\Omega}_{j}},$ (31)
where we introduced $P_{0}:H^{1}(\Omega)\to H^{1}(\Omega)$ defined as: for any $u\in H^{1}(\Omega)$, $P_{0}u\in U_{\tau}$ such that:
$a_{\Omega}(P_{0}u,u_{0}^{t})=a_{\Omega}(u,u_{0}^{t}),\quad\forall u_{0}^{t}\in U_{\tau}.$ (32)
Provided Assumption 1 holds, the operator $P_{0}$ is well-defined. Besides, $P_{0}$ is an oblique projection on the closed space $U_{\tau}$ and the operator norm of $\mathrm{I}-P_{0}$ for $c_{\Omega}$ is bounded by the stability constant $\sigma$ (see (27)) of the coarse problem, namely
$\sup_{0\neq u\in H^{1}(\Omega)}\!\!\frac{\|(\mathrm{I}-P_{0})u\|_{\Omega}}{\|u\|_{\Omega}}\leq\sigma.$ (33)
We can now state our main result, which is the counterpart for the heterogeneous Helmholtz problem of the matrix formulation result stated in (Nataf2024, , Th. 2.12).
Theorem 2.1
If the user-controlled parameter $\tau$ is taken to be sufficiently small such that Assumption 1 holds and $\rho:=\sigma\sqrt{k_{0}k_{1}\tau}<1$, then the iterative solution $u^{n}_{\mathrm{II}}$ defined in (22) converges towards the exact solution $u$ of the model problem (2), in the norm given by $c_{\Omega}$ defined in (13), and with rate $\rho$, namely
$\|u-u^{n}_{\mathrm{II}}\|_{\Omega}\leq\rho^{n}\|u-u^{0}_{\mathrm{II}}\|_{\Omega},\qquad n\geq 0.$ (34)
Proof
Let $\Pi_{j}$ be the $c_{\widetilde{\Omega}_{j}}$-orthogonal projection on the finite dimensional space
[…]
Then, by construction of the coarse space, the image of $\sum_{j}E_{j}\chi_{j}R_{j}\Pi_{j}$ is in the kernel of $\mathrm{I}-P_{0}$, hence
$e^{n+1}_{\mathrm{II}}=(\mathrm{I}-P_{0})\sum_{j}E_{j}\chi_{j}R_{j}[\Pi_{j}+(\mathrm{I}-\Pi_{j})]e^{n}_{\mathrm{II}}|_{\widetilde{\Omega}_{j}}=(\mathrm{I}-P_{0})\sum_{j}E_{j}\chi_{j}R_{j}(\mathrm{I}-\Pi_{j})e^{n}_{\mathrm{II}}|_{\widetilde{\Omega}_{j}},$ (36)
and moreover (see e.g. (Dolean2015, , Lem. 7.7))
$\|\chi_{j}R_{j}(\mathrm{I}-\Pi_{j})e^{n}_{\mathrm{II}}|_{\widetilde{\Omega}_{j}}\|_{\Omega_{j}}^{2}\leq\tau\|(\mathrm{I}-\Pi_{j})e^{n}_{\mathrm{II}}|_{\widetilde{\Omega}_{j}}\|_{\widetilde{\Omega}_{j}}^{2}\leq\tau\|e^{n}_{\mathrm{II}}|_{\widetilde{\Omega}_{j}}\|_{\widetilde{\Omega}_{j}}^{2}.$ (37)
The error can hence be estimated as follows
$\|e^{n+1}_{\mathrm{II}}\|_{\Omega}^{2}$ $\leq\sigma^{2}\Big\|\sum_{j}E_{j}\chi_{j}R_{j}(\mathrm{I}-\Pi_{j})e^{n}_{\mathrm{II}}|_{\widetilde{\Omega}_{j}}\Big\|_{\Omega}^{2}\leq\sigma^{2}k_{0}\sum_{j}\|\chi_{j}R_{j}(\mathrm{I}-\Pi_{j})e^{n}_{\mathrm{II}}|_{\widetilde{\Omega}_{j}}\|_{\Omega_{j}}^{2}$
$\leq\sigma^{2}k_{0}\tau\sum_{j}\|e^{n}_{\mathrm{II}}|_{\widetilde{\Omega}_{j}}\|_{\widetilde{\Omega}_{j}}^{2}\leq\sigma^{2}k_{0}k_{1}\tau\|e^{n}_{\mathrm{II}}\|_{\Omega}^{2}.$ (38)
In practice, Krylov acceleration is used, and one iteration of the above algorithm (22) is used as a preconditioner. Theorem 2.1 also implies convergence of the Krylov iterative method.
3 Numerical illustrations
We conclude with some numerical experiments on the Marmousi model, which represents a realistic geological structure with varying velocity $c$ (see Figure 1) and fits within the model problem (1) with $\nu=c^{-2}$, $\mu=1$. A homogeneous Dirichlet boundary condition is applied on the surface and Robin boundary conditions are imposed on the remaining boundaries. The source term $f$ is a regularized Dirac mass (sharp Gaussian function) centered near the surface (and $g=0$).
The implementation is performed in FreeFEM using the FFDDM framework. The discretization consists of $\mathbb{P}_{2}$ Lagrange finite elements. We perform a weak scaling test using a frequency ramp, so that the number of degrees of freedom in each subdomain remains constant around $\sim 36000$ ($\sim 780$ on the subdomain boundary). We compare results for the one-level ORAS preconditioner and the two-level preconditioners for three coarse space sizes. In practice, instead of using the parameter $\tau$ to filter the eigenvalues, we compute a fixed number of eigenpairs per subdomain given as a percentage of the number of degrees of freedom on the boundary of the subdomain (which is an upper bound on the meaningful eigenvectors to include on the coarse space). This strategy avoids computing too many unnecessary eigenpairs and allows more sensible run time comparisons. The numerical results are reported in Table 1 and in Figure 2.
The results show the poor scalability of the one-level preconditioner for increasing $\omega$. For a large enough coarse space, the second-level preconditioner converges in a much reduced number of iterations, albeit still growing with respect to $\omega$. This is however achieved at the expense of a larger setup time (including factorization of local matrices, computation of eigenvalues, assembly and factorization of the coarse problem). At some point, the reduction in iterations obtained by increasing further the coarse space is offset by the larger computational effort for building the coarse space and solving the coarse problem.
Additional numerical results providing a numerical comparison with alternative methods (including the coarse spaces in Hu2024 ; Ma2024 ) are given in Dolean2025 .
one-level two-level $(7.5\%)$ two-level $(10\%)$ two-level $(12.5\%)$
$\omega$ $J$ NN It Tsetup Tsolver CS It Tsetup Tsolver CS It Tsetup Tsolver CS It Tsetup Tsolver
[…]
Table 1: Angular frequency $\omega$, number of subdomains $J$, number of degrees of freedom $N$, number of GMRES iterations to reach a relative residual of $10^{-8}$ (It), setup time (Tsetup) and solver time (Tsolver) for the one-level preconditioner and two-level preconditioners for three coarse space sizes (CS) given as a percentage of the number of degrees of freedom on the skeleton (X%).
Figure 1: Velocity profile (top) and real part of the solution $u$ for $\omega=60\pi$ (bottom).
Figure 2: GMRES iteration count (left) and total run time (right) with respect to $\omega$.

### v2

@@
 We introduce a decomposition of $\Omega$ into $J\in\mathbb{N}$, $J>0$, overlapping open subdomains $\Omega_{j}$, $j=1,\ldots,J$, associated to a partition-of-unity function $\chi_{j}\in W^{1,\infty}(\Omega_{j})$ such that
-$\overline{\Omega}=\bigcup_{j}\overline{\Omega_{j}},\qquad\chi_{j}\geq 0,\ \chi_{j}|_{\partial\Omega_{j}\setminus\partial\Omega}=0,\qquad\sum_{j}E_{j}\chi_{j}u|_{\Omega_{j}}=u,\ \forall u\in H^{1}(\Omega),$ (5)
+$\overline{\Omega}=\textstyle\bigcup_{j}\overline{\Omega_{j}},\qquad\chi_{j}\geq 0,\ \chi_{j}|_{\partial\Omega_{j}\setminus\partial\Omega}=0,\qquad\textstyle\sum_{j}E_{j}\chi_{j}u|_{\Omega_{j}}=u,\ \forall u\in H^{1}(\Omega),$ (5)
 where $E_{j}$ is an extension-by-continuity operator.
-More precisely, we require that $E_{j}:H^{1}(\Omega_{j})\to H^{1}({\Omega})$ continuously and such that $\operatorname{supp}(E_{j}u_{j})\subset\widetilde{\Omega}_{j}$ for any $u_{j}\in H^{1}(\Omega_{j})$ and for some extended open subdomain $\widetilde{\Omega}_{j}$ such that ${\Omega}_{j}\subsetneq\widetilde{\Omega}_{j}\subset\Omega$.
+More precisely, we require that $E_{j}:H^{1}(\Omega_{j})\to H^{1}({\Omega})$ continuously, $E_{j}u_{j}|_{\Omega_{j}}=u_{j}$ and $\operatorname{supp}(E_{j}u_{j})\subset\widetilde{\Omega}_{j}$ for any $u_{j}\in H^{1}(\Omega_{j})$ and for some extended open subdomain $\widetilde{\Omega}_{j}$ such that ${\Omega}_{j}\subsetneq\widetilde{\Omega}_{j}\subset\Omega$.
 Starting from an initial guess $u^{0}_{\mathrm{I}}\in H^{1}(\Omega)$, the one-level (the index I stands for one-level) Optimized Restricted Additive Schwarz (ORAS) algorithm defines the sequence
-$u^{n+1}_{\mathrm{I}}:=u^{n}_{\mathrm{I}}+\sum_{j}E_{j}\chi_{j}v_{j}^{n+1},\qquad n\geq 0,$ (6)
+$u^{n+1}_{\mathrm{I}}:=u^{n}_{\mathrm{I}}+\textstyle\sum_{j}E_{j}\chi_{j}v_{j}^{n+1},\qquad n\geq 0,$ (6)
 where $v_{j}^{n+1}$ is solution of: Find $v_{j}^{n+1}\in H^{1}(\Omega_{j})$ such that:
-$b_{\Omega_{j}}(v_{j}^{n+1},v_{j}^{t})=l_{\widetilde{\Omega}_{j}}(E_{j}v_{j}^{t})-a_{\widetilde{\Omega}_{j}}(u^{n}_{\mathrm{I}}|_{\widetilde{\Omega}_{j}},E_{j}v_{j}^{t}),\qquad\forall v^{t}_{j}\in H^{1}(\Omega_{j}),$ (7)
+$b_{\Omega_{j}}(v_{j}^{n+1},v_{j}^{t})=l_{\Omega}(E_{j}v_{j}^{t})-a_{\Omega}(u^{n}_{\mathrm{I}},E_{j}v_{j}^{t}),\qquad\forall v^{t}_{j}\in H^{1}(\Omega_{j}),$ (7)
 where we introduced, for any $D\subset\Omega$ and any $v,v^{t}\in H^{1}(D)$:
 $b_{D}(v,v^{t}):=a_{D}(v,v^{t})-\imath\omega(v,v^{t})_{L^{2}(\partial D\setminus\partial\Omega)}.$ (8)
-By similar arguments as for the model problem, the local problem (7) is well-posed.
-One-level error analysis. For $n\geq 0$, the one-level error is defined as
-$e^{n+1}_{\mathrm{I}}:=u-u^{n+1}_{\mathrm{I}}=e^{n}_{\mathrm{I}}-\sum_{j}E_{j}\chi_{j}v_{j}^{n+1}=\sum_{j}E_{j}\chi_{j}(e^{n}_{\mathrm{I}}|_{\Omega_{j}}-v_{j}^{n+1}).$ (9)
+By similar arguments as for the model problem, the local problem (7) is well-posed, and we remark that the right-hand-side in (7) can be computed locally in $\widetilde{\Omega}_{j}$ so that $v_{j}^{n+1}$ is equivalently solution of
+$b_{\Omega_{j}}(v_{j}^{n+1},v_{j}^{t})=l_{\widetilde{\Omega}_{j}}((E_{j}v_{j}^{t})|_{\widetilde{\Omega}_{j}})-a_{\widetilde{\Omega}_{j}}(u^{n}_{\mathrm{I}}|_{\widetilde{\Omega}_{j}},(E_{j}v_{j}^{t})|_{\widetilde{\Omega}_{j}}),\qquad\forall v^{t}_{j}\in H^{1}(\Omega_{j}).$ (9)
+For general partitions (where at least three subdomains overlap in some part of the domain $\Omega$), the sequence of $u_{\mathrm{I}}^{n}$ depends a priori on the choice of $E_{j}$ (which is not unique), since it affects the right-hand-side of (7) (in (6), the presence of the partition-of-unity $\chi_{j}$ implies that $E_{j}$ extends by zero, which is uniquely defined).
+One-level error analysis. For $n\geq 0$, the one-level error $e^{n}_{\mathrm{I}}:=u-u^{n}_{\mathrm{I}}$ satisfies
+$e^{n+1}_{\mathrm{I}}=u-u^{n+1}_{\mathrm{I}}=e^{n}_{\mathrm{I}}-\textstyle\sum_{j}E_{j}\chi_{j}v_{j}^{n+1}=\textstyle\sum_{j}E_{j}\chi_{j}(e^{n}_{\mathrm{I}}|_{\Omega_{j}}-v_{j}^{n+1}).$ (10)
 From (9), we have,
-$b_{\Omega_{j}}(e^{n}_{\mathrm{I}}|_{\Omega_{j}}-v_{j}^{n+1},u_{j}^{t})=b_{\Omega_{j}}(e^{n}_{\mathrm{I}}|_{\Omega_{j}},v_{j}^{t})-a_{\widetilde{\Omega}_{j}}(e^{n}_{\mathrm{I}}|_{\widetilde{\Omega}_{j}},E_{j}v_{j}^{t}),\ \forall v_{j}^{t}\in H^{1}(\Omega_{j}),$ (10)
+$b_{\Omega_{j}}(e^{n}_{\mathrm{I}}|_{\Omega_{j}}-v_{j}^{n+1},v_{j}^{t})=b_{\Omega_{j}}(e^{n}_{\mathrm{I}}|_{\Omega_{j}},v_{j}^{t})-a_{\widetilde{\Omega}_{j}}(e^{n}_{\mathrm{I}}|_{\widetilde{\Omega}_{j}},(E_{j}v_{j}^{t})|_{\widetilde{\Omega}_{j}}),\ \forall v_{j}^{t}\in H^{1}(\Omega_{j}),$ (11)
 so that
-$e^{n+1}_{\mathrm{I}}=\sum_{j}E_{j}\chi_{j}R_{j}e^{n}_{\mathrm{I}}|_{\widetilde{\Omega}_{j}},$ (11)
+$e^{n+1}_{\mathrm{I}}=\textstyle\sum_{j}E_{j}\chi_{j}R_{j}e^{n}_{\mathrm{I}}|_{\widetilde{\Omega}_{j}},$ (12)
 where $R_{j}:H^{1}(\widetilde{\Omega}_{j})\to H^{1}(\Omega_{j})$ is defined for any $\widetilde{v}_{j}\in H^{1}(\widetilde{\Omega}_{j})$ as
-$b_{\Omega_{j}}(R_{j}\widetilde{v}_{j},v_{j}^{t})=b_{\Omega_{j}}(\widetilde{v}_{j}|_{\Omega_{j}},v_{j}^{t})-a_{\widetilde{\Omega}_{j}}(\widetilde{v}_{j},E_{j}v_{j}^{t}),\ \forall v_{j}^{t}\in H^{1}(\Omega_{j}).$ (12)
+$b_{\Omega_{j}}(R_{j}\widetilde{v}_{j},v_{j}^{t})=b_{\Omega_{j}}(\widetilde{v}_{j}|_{\Omega_{j}},v_{j}^{t})-a_{\widetilde{\Omega}_{j}}(\widetilde{v}_{j},(E_{j}v_{j}^{t})|_{\widetilde{\Omega}_{j}}),\qquad\forall v_{j}^{t}\in H^{1}(\Omega_{j}).$ (13)
 The operator $R_{j}$ is well-defined since the local problems (13) are well-posed.
@@
 The error can be estimated as follows
-$\|e^{n+1}_{\mathrm{I}}\|_{\Omega}^{2}\leq k_{0}\sum_{j}\|\chi_{j}R_{j}e^{n}_{\mathrm{I}}|_{\widetilde{\Omega}_{j}}\|_{\Omega_{j}}^{2}\leq k_{0}\xi\sum_{j}\|e^{n}_{\mathrm{I}}|_{\widetilde{\Omega}_{j}}\|_{\widetilde{\Omega}_{j}}^{2}\leq k_{0}k_{1}\xi\|e^{n}_{\mathrm{I}}\|_{\Omega}^{2},$ (14)
+$\|e^{n+1}_{\mathrm{I}}\|_{\Omega}^{2}\leq k_{0}\textstyle\sum_{j}\big\|\chi_{j}R_{j}e^{n}_{\mathrm{I}}|_{\widetilde{\Omega}_{j}}\big\|_{\Omega_{j}}^{2}\leq k_{0}\xi\textstyle\sum_{j}\big\|e^{n}_{\mathrm{I}}|_{\widetilde{\Omega}_{j}}\big\|_{\widetilde{\Omega}_{j}}^{2}\leq k_{0}k_{1}\xi\|e^{n}_{\mathrm{I}}\|_{\Omega}^{2},$ (15)
 where we introduced
-$k_{0}:=\!\!\sup_{\begin{subarray}{c}(u_{j})_{j}\\ 0\neq u_{j}\in H^{1}_{0}(\Omega_{j})\end{subarray}}\!\!\frac{\|\sum_{j}E_{j}u_{j}\|_{\Omega}^{2}}{\sum_{j}\|u_{j}\|_{\Omega_{j}}^{2}},\qquad k_{1}:=\!\!\sup_{0\neq u\in H^{1}(\Omega)}\!\!\frac{\sum_{j}\|u|_{\widetilde{\Omega}_{j}}\|_{\widetilde{\Omega}_{j}}^{2}}{\|u\|_{\Omega}^{2}},$ (15)
+$k_{0}:=\!\!\sup_{\begin{subarray}{c}(u_{j})_{j}\\ 0\neq u_{j}\in H^{1}_{0}(\Omega_{j})\end{subarray}}\!\!\frac{\|\textstyle\sum_{j}E_{j}u_{j}\|_{\Omega}^{2}}{\textstyle\sum_{j}\|u_{j}\|_{\Omega_{j}}^{2}},\qquad k_{1}:=\!\!\sup_{0\neq u\in H^{1}(\Omega)}\!\!\frac{\textstyle\sum_{j}\|u|_{\widetilde{\Omega}_{j}}\|_{\widetilde{\Omega}_{j}}^{2}}{\|u\|_{\Omega}^{2}},$ (16)
 $\xi:=\sup_{j}\xi_{j},\qquad\xi_{j}:=\!\!\sup_{0\neq\widetilde{u}_{j}\in H^{1}(\widetilde{\Omega}_{j})}\!\!\frac{\|\chi_{j}R_{j}\widetilde{u}_{j}\|_{\Omega_{j}}^{2}}{\|\widetilde{u}_{j}\|_{\widetilde{\Omega}_{j}}^{2}}.$ (17)
@@
 The two constants $k_{0}$ and $k_{1}$ can be bounded uniformly of the parameters $\mu,\nu,\omega$ of the problem and only depend respectively on the maximum number of neighbors for a subdomain (see e.g.
-(Dolean2015, , Lem.
+(Dolean2015, Lem.
 7.9)) and the maximum number of subdomain intersection multiplicity (see e.g.
-(Dolean2015, , Lem.
+(Dolean2015, Lem.
 7.13)).
@@
 Find $(\widetilde{u}_{j},v_{j},\sigma_{j},\lambda_{j})\in H^{1}(\widetilde{\Omega}_{j})\times H^{1}(\Omega_{j})\times H^{1}(\Omega_{j})\times\mathbb{R}$ such that:
-$\begin{cases}c_{\Omega_{j}}(\chi_{j}v_{j},\chi_{j}v_{j}^{t})-b_{\Omega_{j}}(v_{j}^{t},\sigma_{j})=0,&\quad\forall v_{j}^{t}\in H^{1}(\Omega_{j}),\\ -b_{\Omega_{j}}(v_{j},\sigma_{j}^{t})+b_{\Omega_{j}}(\widetilde{u}_{j}|_{\Omega_{j}},\sigma_{j}^{t})-a_{\widetilde{\Omega}_{j}}(\widetilde{u}_{j},E_{j}\sigma_{j}^{t})=0,&\quad\forall\sigma_{j}^{t}\in H^{1}(\Omega_{j}),\\ b_{\Omega_{j}}(\widetilde{u}_{j}^{t}|_{\Omega_{j}},\sigma_{j})-a_{\widetilde{\Omega}_{j}}(\widetilde{u}_{j}^{t},E_{j}\sigma_{j})=\lambda_{j}c_{\widetilde{\Omega}_{j}}(\widetilde{u}_{j},\widetilde{u}_{j}^{t}),&\quad\forall\widetilde{u}_{j}^{t}\in H^{1}(\widetilde{\Omega}_{j}).\end{cases}$ (18)
+$\begin{cases}c_{\Omega_{j}}(\chi_{j}v_{j},\chi_{j}v_{j}^{t})-b_{\Omega_{j}}(v_{j}^{t},\sigma_{j})=0,&\forall v_{j}^{t}\in H^{1}(\Omega_{j}),\\ -b_{\Omega_{j}}(v_{j},\sigma_{j}^{t})+b_{\Omega_{j}}(\widetilde{u}_{j}|_{\Omega_{j}},\sigma_{j}^{t})-a_{\widetilde{\Omega}_{j}}(\widetilde{u}_{j},(E_{j}\sigma_{j}^{t})|_{\widetilde{\Omega}_{j}})=0,&\forall\sigma_{j}^{t}\in H^{1}(\Omega_{j}),\\ b_{\Omega_{j}}(\widetilde{u}_{j}^{t}|_{\Omega_{j}},\sigma_{j})-a_{\widetilde{\Omega}_{j}}(\widetilde{u}_{j}^{t},(E_{j}\sigma_{j})|_{\widetilde{\Omega}_{j}})=\lambda_{j}c_{\widetilde{\Omega}_{j}}(\widetilde{u}_{j},\widetilde{u}_{j}^{t}),&\forall\widetilde{u}_{j}^{t}\in H^{1}(\widetilde{\Omega}_{j}).\end{cases}$ (19)
 The following result shows that each $\xi_{j}$ (hence $\xi$) is finite and that the eigenvalues $\lambda_{j}$ solutions to (18) converge to 0.
@@
 Proof
-The result is based on a Caccioppoli inequality which we establish next (see also (Ma2023, , Lem. 2.7)). From (12) we have, for any $\widetilde{u}_{j}\in H^{1}(\widetilde{\Omega}_{j})$,
+The result is based on a Caccioppoli inequality which we establish next (see also (Ma2023, Lem. 2.7)). From (13) we have, for any $\widetilde{u}_{j}\in H^{1}(\widetilde{\Omega}_{j})$,
 $b_{\Omega_{j}}(R_{j}\widetilde{u}_{j},\chi_{j}^{2}R_{j}\widetilde{u}_{j})=0,$ (20)
@@
 $(\mu\chi_{j}\nabla R_{j}\widetilde{u}_{j},\chi_{j}\nabla R_{j}\widetilde{u}_{j})_{L^{2}(\Omega_{j})^{d}}+(\mu R_{j}\widetilde{u}_{j}\nabla\chi_{j},\chi_{j}\nabla R_{j}\widetilde{u}_{j})_{L^{2}(\Omega_{j})^{d}}$
-$\quad+(\mu\chi_{j}\nabla R_{j}\widetilde{u}_{j},R_{j}\widetilde{u}_{j}\nabla\chi_{j})_{L^{2}(\Omega_{j})^{d}}-\omega^{2}(\nu R_{j}\widetilde{u}_{j},\chi_{j}^{2}R_{j}\widetilde{u}_{j})_{L^{2}(\Omega_{j})}=0.$
+$\quad+(\mu\chi_{j}\nabla R_{j}\widetilde{u}_{j},R_{j}\widetilde{u}_{j}\nabla\chi_{j})_{L^{2}(\Omega_{j})^{d}}-\omega^{2}(\nu R_{j}\widetilde{u}_{j},\chi_{j}^{2}R_{j}\widetilde{u}_{j})_{L^{2}(\Omega_{j})}=0.$ (21)
 On the other hand, we have, for any $\widetilde{u}_{j}\in H^{1}(\widetilde{\Omega}_{j})$,
@@
 Starting from an initial guess $u^{0}_{\mathrm{II}}\in H^{1}(\Omega)$, we define a sequence of solutions $u^{n}_{\mathrm{II}}\in H^{1}(\Omega)$ (the index II stands for two-level),
-$u^{n+1}_{\mathrm{II}}:=u^{n}_{\mathrm{II}}+\sum_{j}E_{j}\chi_{j}w_{j}^{n+1}+w_{0}^{n+1},\qquad n\geq 0,$ (22)
+$u^{n+1}_{\mathrm{II}}:=u^{n}_{\mathrm{II}}+\textstyle\sum_{j}E_{j}\chi_{j}w_{j}^{n+1}+w_{0}^{n+1},\qquad n\geq 0,$ (24)
 where $w_{j}^{n+1}$ is solution of: Find $w_{j}^{n+1}\in H^{1}(\Omega_{j})$ such that:
-$b_{\Omega_{j}}(w_{j}^{n+1},w_{j}^{t})=l_{\widetilde{\Omega}_{j}}(E_{j}w_{j}^{t})-a_{\widetilde{\Omega}_{j}}(u^{n}_{\mathrm{II}}|_{\widetilde{\Omega}_{j}},E_{j}w_{j}^{t}),\qquad\forall w^{t}_{j}\in H^{1}(\Omega_{j}),$ (23)
+$b_{\Omega_{j}}(w_{j}^{n+1},w_{j}^{t})=l_{\Omega}(E_{j}w_{j}^{t})-a_{\Omega}(u^{n}_{\mathrm{II}},E_{j}w_{j}^{t}),\qquad\forall w^{t}_{j}\in H^{1}(\Omega_{j}),$ (25)
 and $w_{0}^{n+1}$ is solution of: Find $w_{0}^{n+1}\in U_{\tau}$ such that:
-$a_{\Omega}(w_{0}^{n+1},w_{0}^{t})=l_{\Omega}(w_{0}^{t})-a_{\Omega}(u^{n}_{\mathrm{II}}+\sum_{j}E_{j}\chi_{j}w_{j}^{n+1},w_{0}^{t}),\quad\forall w_{0}^{t}\in U_{\tau}.$ (24)
+$a_{\Omega}(w_{0}^{n+1},w_{0}^{t})=l_{\Omega}(w_{0}^{t})-a_{\Omega}(u^{n}_{\mathrm{II}}+\textstyle\sum_{j}E_{j}\chi_{j}w_{j}^{n+1},w_{0}^{t}),\quad\forall w_{0}^{t}\in U_{\tau}.$ (26)
 Assumption 1
@@
 taking the user-controlled parameter $\tau$ sufficiently small).
-For instance, well-posedness of the coarse problem and a quasi-optimality estimate implying (27) can be obtained using the so-called Aubin–Nitsche duality trick and Schatz argument, see (Galkowski2025, , Lemma 5.1).
-Two-level error analysis. For $n\geq 0$, the two-level error is defined as
-$e^{n+1}_{\mathrm{II}}:=u-u^{n+1}_{\mathrm{II}}=u-u^{n}_{\mathrm{II}}-\sum_{j}E_{j}\chi_{j}w_{j}^{n+1}-w_{0}^{n+1}.$ (28)
+For instance, well-posedness of the coarse problem and a quasi-optimality estimate implying (29) can be obtained using the so-called Aubin–Nitsche duality trick and Schatz argument, see (Galkowski2025, Lemma 5.1).
+Two-level error analysis. For $n\geq 0$, the two-level error $e^{n}_{\mathrm{II}}:=u-u^{n}_{\mathrm{II}}$ satisfies
+$e^{n+1}_{\mathrm{II}}=u-u^{n+1}_{\mathrm{II}}=u-u^{n}_{\mathrm{II}}-\textstyle\sum_{j}E_{j}\chi_{j}w_{j}^{n+1}-w_{0}^{n+1}.$ (30)
 From (25), we have,
-$b_{\Omega_{j}}(e^{n}_{\mathrm{II}}|_{\Omega_{j}}-w_{j}^{n+1},w_{j}^{t})=b_{\Omega_{j}}(e^{n}_{\mathrm{II}}|_{\Omega_{j}},w_{j}^{t})-a_{\widetilde{\Omega}_{j}}(e^{n}_{\mathrm{II}}|_{\widetilde{\Omega}_{j}},E_{j}w_{j}^{t}),\ \forall w_{j}^{t}\in H^{1}(\Omega_{j}),$ (29)
+$b_{\Omega_{j}}(e^{n}_{\mathrm{II}}|_{\Omega_{j}}-w_{j}^{n+1},w_{j}^{t})=b_{\Omega_{j}}(e^{n}_{\mathrm{II}}|_{\Omega_{j}},w_{j}^{t})-a_{\widetilde{\Omega}_{j}}(e^{n}_{\mathrm{II}}|_{\widetilde{\Omega}_{j}},(E_{j}w_{j}^{t})|_{\widetilde{\Omega}_{j}}),\ \forall w_{j}^{t}\in H^{1}(\Omega_{j}),$ (31)
 and from (26), we have
-$a_{\Omega}(w_{0}^{n+1},w_{0}^{t})=a_{\Omega}(\sum_{j}E_{j}\chi_{j}(e^{n}_{\mathrm{II}}|_{\Omega_{j}}-w_{j}^{n+1}),w_{0}^{t}),\quad\forall w_{0}^{t}\in U_{\tau},$ (30)
+$a_{\Omega}(w_{0}^{n+1},w_{0}^{t})=a_{\Omega}(\textstyle\sum_{j}E_{j}\chi_{j}(e^{n}_{\mathrm{II}}|_{\Omega_{j}}-w_{j}^{n+1}),w_{0}^{t}),\quad\forall w_{0}^{t}\in U_{\tau},$ (32)
 so that
-$e^{n+1}_{\mathrm{II}}=(\mathrm{I}-P_{0})\sum_{j}E_{j}\chi_{j}R_{j}e^{n}_{\mathrm{II}}|_{\widetilde{\Omega}_{j}},$ (31)
+$e^{n+1}_{\mathrm{II}}=(\mathrm{I}-P_{0})\textstyle\sum_{j}E_{j}\chi_{j}R_{j}e^{n}_{\mathrm{II}}|_{\widetilde{\Omega}_{j}},$ (33)
 where we introduced $P_{0}:H^{1}(\Omega)\to H^{1}(\Omega)$ defined as: for any $u\in H^{1}(\Omega)$, $P_{0}u\in U_{\tau}$ such that:
@@
 $\sup_{0\neq u\in H^{1}(\Omega)}\!\!\frac{\|(\mathrm{I}-P_{0})u\|_{\Omega}}{\|u\|_{\Omega}}\leq\sigma.$ (35)
-We can now state our main result, which is the counterpart for the heterogeneous Helmholtz problem of the matrix formulation result stated in (Nataf2024, , Th. 2.12).
+We can now state our main result, which is the counterpart for the heterogeneous Helmholtz problem of the matrix formulation result stated in (Nataf2025, Th. 2.12).
 Theorem 2.1
@@
 Then, by construction of the coarse space, the image of $\sum_{j}E_{j}\chi_{j}R_{j}\Pi_{j}$ is in the kernel of $\mathrm{I}-P_{0}$, hence
-$e^{n+1}_{\mathrm{II}}=(\mathrm{I}-P_{0})\sum_{j}E_{j}\chi_{j}R_{j}[\Pi_{j}+(\mathrm{I}-\Pi_{j})]e^{n}_{\mathrm{II}}|_{\widetilde{\Omega}_{j}}=(\mathrm{I}-P_{0})\sum_{j}E_{j}\chi_{j}R_{j}(\mathrm{I}-\Pi_{j})e^{n}_{\mathrm{II}}|_{\widetilde{\Omega}_{j}},$ (36)
-and moreover (see e.g. (Dolean2015, , Lem. 7.7))
+$e^{n+1}_{\mathrm{II}}=(\mathrm{I}-P_{0})\textstyle\sum_{j}E_{j}\chi_{j}R_{j}[\Pi_{j}+(\mathrm{I}-\Pi_{j})]e^{n}_{\mathrm{II}}|_{\widetilde{\Omega}_{j}}=(\mathrm{I}-P_{0})\textstyle\sum_{j}E_{j}\chi_{j}R_{j}(\mathrm{I}-\Pi_{j})e^{n}_{\mathrm{II}}|_{\widetilde{\Omega}_{j}},$ (38)
+and moreover (see e.g. (Dolean2015, Lem. 7.7))
 $\|\chi_{j}R_{j}(\mathrm{I}-\Pi_{j})e^{n}_{\mathrm{II}}|_{\widetilde{\Omega}_{j}}\|_{\Omega_{j}}^{2}\leq\tau\|(\mathrm{I}-\Pi_{j})e^{n}_{\mathrm{II}}|_{\widetilde{\Omega}_{j}}\|_{\widetilde{\Omega}_{j}}^{2}\leq\tau\|e^{n}_{\mathrm{II}}|_{\widetilde{\Omega}_{j}}\|_{\widetilde{\Omega}_{j}}^{2}.$ (39)
-The error can hence be estimated as follows
-$\|e^{n+1}_{\mathrm{II}}\|_{\Omega}^{2}$ $\leq\sigma^{2}\Big\|\sum_{j}E_{j}\chi_{j}R_{j}(\mathrm{I}-\Pi_{j})e^{n}_{\mathrm{II}}|_{\widetilde{\Omega}_{j}}\Big\|_{\Omega}^{2}\leq\sigma^{2}k_{0}\sum_{j}\|\chi_{j}R_{j}(\mathrm{I}-\Pi_{j})e^{n}_{\mathrm{II}}|_{\widetilde{\Omega}_{j}}\|_{\Omega_{j}}^{2}$
-$\leq\sigma^{2}k_{0}\tau\sum_{j}\|e^{n}_{\mathrm{II}}|_{\widetilde{\Omega}_{j}}\|_{\widetilde{\Omega}_{j}}^{2}\leq\sigma^{2}k_{0}k_{1}\tau\|e^{n}_{\mathrm{II}}\|_{\Omega}^{2}.$ (38)
-In practice, Krylov acceleration is used, and one iteration of the above algorithm (22) is used as a preconditioner. Theorem 2.1 also implies convergence of the Krylov iterative method.
+Recalling (16) and (35), the error can hence be estimated as follows
+$\|e^{n+1}_{\mathrm{II}}\|_{\Omega}^{2}$ $=\big\|(\mathrm{I}-P_{0})\textstyle\sum_{j}E_{j}\chi_{j}R_{j}(\mathrm{I}-\Pi_{j})e^{n}_{\mathrm{II}}|_{\widetilde{\Omega}_{j}}\big\|_{\Omega}^{2}$
+$\leq\sigma^{2}\;\big\|\textstyle\sum_{j}E_{j}\chi_{j}R_{j}(\mathrm{I}-\Pi_{j})e^{n}_{\mathrm{II}}|_{\widetilde{\Omega}_{j}}\big\|_{\Omega}^{2}\leq\sigma^{2}k_{0}\;\ \textstyle\sum_{j}\big\|\chi_{j}R_{j}(\mathrm{I}-\Pi_{j})e^{n}_{\mathrm{II}}|_{\widetilde{\Omega}_{j}}\big\|_{\Omega_{j}}^{2}$
+$\leq\sigma^{2}k_{0}\tau\;\textstyle\sum_{j}\big\|e^{n}_{\mathrm{II}}|_{\widetilde{\Omega}_{j}}\big\|_{\widetilde{\Omega}_{j}}^{2}\leq\sigma^{2}k_{0}k_{1}\tau\;\|e^{n}_{\mathrm{II}}\|_{\Omega}^{2}.$ (40)
+In practice, Krylov acceleration is used, and one iteration of the above algorithm (24) is used as a preconditioner.
+Theorem 2.1 also implies convergence of the Krylov iterative method with this preconditioner.
 3 Numerical illustrations
@@
 At some point, the reduction in iterations obtained by increasing further the coarse space is offset by the larger computational effort for building the coarse space and solving the coarse problem.
-Additional numerical results providing a numerical comparison with alternative methods (including the coarse spaces in Hu2024 ; Ma2024 ) are given in Dolean2025 .
+Additional numerical results providing a numerical comparison with alternative methods (including the coarse spaces in Hu2025; Ma2025b) are given in Dolean2025.
 one-level two-level $(7.5\%)$ two-level $(10\%)$ two-level $(12.5\%)$